# biblio.bib

### v2

@@ Abstract.
Nanodevices with Terahertz (THz)-based wireless communication capabilities are providing a primer for flow-guided localization within the human bloodstreams. Such localization is allowing for assigning the locations of sensed events with the events themselves, providing benefits in precision medicine along the lines of early and precise diagnostics, and reduced costs and invasiveness. Flow-guided localization is still in a rudimentary phase, with only a handful of works targeting the problem. Nonetheless, the performance assessments of the proposed solutions are already carried out in a non-standardized way, usually along a single performance metric, and ignoring various aspects that are relevant at such a scale (e.g., nanodevices’ limited energy) and for such a challenging environment (e.g., extreme attenuation of in-body THz propagation). As such, these assessments feature low levels of realism and cannot be compared in an objective way. Toward addressing this issue, we account for the environmental and scale-related peculiarities of the scenario and assess the performance of two state-of-the-art flow-guided localization approaches along a set of heterogeneous performance metrics such as the accuracy and reliability of localization.
Machine Learning, Terahertz, Graph Neural Networks, Radio Frequency, Neural Network, Nanodevice, Localization
conference: Woodstock ’18: ACM Symposium on Neural Gaze Detection; June 03–05, 2018; Woodstock, NYccs: Networks Network architecturesccs: Networks Network simulationsccs: Networks Location based servicesccs: Human-centered computing Ubiquitous and mobile computing
@@ 1. Introduction
In this article, we aim to enhance such assessments’ realism by simultaneously considering multiple factors. These factors include accounting for the nanodevices’ mobility, in-body nanoscale THz communication between them and the external world, and various energy-related and technological constraints, such as pulse-based modulation, which impact the nanodevices’ performance. By incorporating these elements through the utilization of a simulator for objective and standardized performance evaluation of flow-guided nanoscale localization (lopez2023toward), we seek to provide a more comprehensive and realistic understanding of the performance of flow-guided localization. To the best of our knowledge, this article represents one of the first attempts at an objective and realistic performance assessment of different flow-guided localization approaches across a set of standardized heterogeneous metrics, including point and region estimation accuracies and localization reliability. Our approach is inspired by established approaches targeting objective evaluation and benchmarking of more traditional indoor localization approaches stemming from efforts such as NIST PerfLoc (moayeri2016perfloc), EU EVARILOS (van2015platform), and Microsoft/IPSN indoor localization competition (lymberopoulos2015realistic).
@@ 2.1. Flow-guided Localization Fundamentals
Flow-guided nanoscale localization aims to detect and locate a target event using nanodevices without requiring the nanodevices to determine their own location. The concept presented in (lemic2022toward) supports this type of scenario and falls within this category. However, the notable representatives of this localization approach are (gomez2022nanosensor) and (simonjan2021body). In these studies, Machine Learning (ML) models are employed to differentiate the regions through which each nanodevice passes during a single circulation through the bloodstream. The authors in (simonjan2021body) achieve this by tracking the distances a nanodevice covers during its circulation using a nanoscale Inertial Measurement Unit (IMU). However, this poses challenges concerning the limited resources available for storing and processing IMU-generated data at the nanodevice level, as well as the accuracy of IMU readings being affected by the blood’s vortex flow. On the other hand, the authors in (gomez2022nanosensor) address these challenges by tracking the time required for each circulation through the bloodstream. The captured distance or time information is then transmitted to a beaconing anchor located near the heart using short-range THz-based backscattering at the nanodevice level.
@@ 4. Evaluation Results
Figure 2 depicts the performance achieved by the two solutions as a function of the number of administered nanodevices. The figure shows that the in-house solution generally outperforms the off-the-shelf solution regarding the point accuracy metric. For example, in the scenario assuming 128 nanodevices deployed within the bloodstream, the in-house solution yields a median localization error of around 25 cm throughout the simulation run. In contrast, this error is roughly 40 cm throughout the simulation for the off-the-shelf solution. This can be confirmed by the results depicted in Figures 3 and 4 demonstrating the influence of the event sampling granularity and event detection threshold of the performance of the considered solutions, respectively. A similar conclusion can be drawn when observing the region detection accuracy, although in this case, the improvements yielded by the in-house solution are less pronounced and in the range of several percent. This behavior is observed because the off-the-shelf solution is not tailored to our framework data. As such, its hyperparameters have not been fine-tuned for the raw datasets utilized in this study, in contrast to the in-house-made and finely-tuned solution.
@@
Regarding the design space of the considered solutions, Figure 2 depicts the effects of an increasing number of administered nanodevices on the accuracy and reliability-related metrics. As visible, an increase in the number of nanodevices increases the accuracies of both solutions for point and region estimation. As an example, when increasing the number of nanodevices from 32 to 128 and considering an entire simulation run with the duration of 960 sec, the third quartile of localization error for the in-house solution is decreased from more than 75 to less than 40 cm. This improvement is a direct result of the system producing more raw data as a result of introducing additional nanodevices. However, the region accuracy is only minorly affected, which can be attributed to the fact that the more data is generated, the more erroneous instances it features, eventually resulting in poor region detection accuracy. The introduction of additional nanodevices benefits reliability, although eventually, the reliability of producing location estimates across solutions and scenarios converges toward 100%. Optimizing the reliability increase will interest applications with latency requirements, e.g., targeted drug delivery.
Figures 3 and 4 depict the effects of the event sampling frequency and detection distance threshold on the performance of the solutions, respectively. As visible from the figures, these parameters’ variations only slightly affect the considered metrics. The sampling granularity only affects the rate at which the events can be detected. As such, it is expected that the more frequent the sampling, the higher the convergence rate toward the region detection accuracy achievable by the models. In addition, the energy consumed in sampling is small compared to the overall energy of the nanodevices. In other words, even the fastest sampling of 10 samples/sec is insufficient to deplete the energy of the nanodevices; hence, the granularity does not significantly affect the reliability of producing estimates. In more general terms, the selection of sampling frequency should be considered in relation to the blood speeds and available energy at the nanonode level. For the considered regions in which the blood speeds are 1 cm/sec and with the event detection threshold of 1 cm, there is no need for increasing the sampling frequency beyond 1 sample/sec, as this will solely result in increased energy consumption.

### v3

# Insights from the Design Space Exploration of Flow-Guided Nanoscale Localization

@@ Abstract.
Nanodevices with Terahertz (THz)-based wireless communication capabilities are providing a primer for flow-guided localization within the human bloodstreams. Such localization is allowing for assigning the locations of sensed events with the events themselves, providing benefits along the lines of early and precise diagnostics, and reduced costs and invasiveness. Flow-guided localization is still in a rudimentary phase, with only a handful of works targeting the problem. Nonetheless, the performance assessments of the proposed solutions are already carried out in a non-standardized way, usually along a single performance metric, and ignoring various aspects that are relevant at such a scale (e.g., nanodevices’ limited energy) and for such a challenging environment (e.g., extreme attenuation of in-body THz propagation). As such, these assessments feature low levels of realism and cannot be compared in an objective way. Toward addressing this issue, we account for the environmental and scale-related peculiarities of the scenario and assess the performance of two state-of-the-art flow-guided localization approaches along a set of heterogeneous performance metrics such as the accuracy and reliability of localization.
ccs: Networks Network architecturesccs: Networks Network simulationsccs: Networks Location based servicesccs: Human-centered computing
@@ 1. Introduction
In this article, we aim to enhance such assessments’ realism by simultaneously considering multiple factors. These factors include accounting for the nanodevices’ mobility, in-body nanoscale THz communication between them and the external world, and various energy-related and technological constraints, such as pulse-based modulation, which impact the nanodevices’ performance. By incorporating these elements through the utilization of a simulator for objective and standardized performance evaluation of flow-guided nanoscale localization (lopez2023toward), we seek to provide a more comprehensive and realistic understanding of the performance of flow-guided localization. To the best of our knowledge, this article represents one of the first attempts at an objective and realistic performance assessment of different flow-guided localization approaches across a set of standardized heterogeneous metrics, including point and region estimation accuracies and localization reliability. Our work is inspired by established approaches targeting objective evaluation and benchmarking of traditional indoor localization solutions stemming from efforts such as NIST PerfLoc (moayeri2016perfloc), EU EVARILOS (van2015platform), and Microsoft/IPSN indoor localization competition (lymberopoulos2015realistic).
@@ 2.1. Flow-guided Localization Fundamentals
Flow-guided localization aims to localize a target event using nanodevices without requiring the nanodevices to determine their own location. The concept presented in (lemic2022toward) supports this type of scenario and falls within this category. However, the notable representatives of this localization approach are (gomez2022nanosensor) and (simonjan2021body). In these studies, Machine Learning (ML) models are employed to differentiate the regions through which each nanodevice passes during a single circulation through the bloodstream. The authors in (simonjan2021body) achieve this by tracking the distances a nanodevice covers during its circulation using a nanoscale Inertial Measurement Unit (IMU). However, this poses challenges concerning the limited resources available for storing and processing IMU-generated data at the nanodevice level, as well as the accuracy of IMU readings being affected by the blood’s vortex flow. On the other hand, the authors in (gomez2022nanosensor) address these challenges by tracking the time required for each circulation through the bloodstream. The captured distance or time information is then transmitted to a beaconing anchor located near the heart using short-range THz-based backscattering.
@@ 4. Evaluation Results
Figure 2 depicts the performance achieved by the two solutions as a function of the number of administered nanodevices. The figure shows that the in-house solution generally outperforms the off-the-shelf solution regarding the point accuracy metric. For example, in the scenario assuming 128 deployed nanodevices, the in-house solution yields a median localization error of around 25 cm throughout the simulation run. In contrast, this error is roughly 40 cm throughout the simulation for the off-the-shelf solution. This can be confirmed by the results depicted in Figures 3 and 4 demonstrating the influence of the event sampling granularity and event detection threshold of the performance of the considered solutions, respectively. A similar conclusion can be drawn when observing the region detection accuracy, although in this case, the improvements yielded by the in-house solution are less pronounced and in the range of several percent. This is observed because the off-the-shelf solution is not tailored to our framework data. As such, its hyperparameters have not been fine-tuned for the raw datasets utilized in this study, in contrast to the in-house-made solution.
@@
Regarding the design space of the considered solutions, Figure 2 depicts the effects of an increasing number of administered nanodevices on the accuracy and reliability-related metrics. As visible, an increase in the number of nanodevices increases the accuracies of both solutions for point and region estimation. As an example, when increasing the number of nanodevices from 32 to 128 and considering an entire simulation run with the duration of 960 sec, the third quartile of localization error for the in-house solution is decreased from more than 75 to less than 40 cm. This improvement is a direct result of the system producing more raw data as a result of introducing additional nanodevices. However, the region accuracy is only minorly affected, which can be attributed to the fact that the more data is generated, the more erroneous instances it features, eventually resulting in poor region detection accuracy. The introduction of additional nanodevices benefits reliability, although eventually, the reliability of producing location estimates generally converges toward 100%. Optimizing the reliability increase will benefit applications with latency requirements, e.g., drug delivery.
@@
Figures 3 and 4 depict the effects of the event sampling frequency and detection distance threshold on the performance of the solutions, respectively. As visible, these parameters’ variations only slightly affect the considered metrics. The sampling granularity only affects the rate at which the events can be detected. As such, it is expected that the more frequent the sampling, the higher the convergence rate toward the region detection accuracy achievable by the models. In addition, the energy consumed in sampling is small compared to the overall energy of the nanodevices. In other words, even the fastest sampling of 10 samples/sec is insufficient to deplete the energy of the nanodevices; hence, the granularity does not significantly affect the reliability of producing estimates. More generally, the selection of sampling frequency should be considered in relation to the blood speeds and available energy at the nanonode level. For the considered regions in which the blood speeds are 1 cm/sec and with the event detection threshold of 1 cm, there is no need for increasing the sampling frequency beyond 1 sample/sec, as this will solely result in increased energy consumption.
@@ 5. Conclusion
Acknowledgments
This work was supported in part by the project IoBNT, funded by BMBF under grant number 16KIS1986K, and by the project NaBoCom, funded by DFG under grant number DR 639/21-2.